OpenStudy (anonymous):

.... f(x)=x^2+x and g(x)=2x+1 solve for (f o g)(2)

hero (hero):

f o g (2) = f(g(2)) = (5)^2 + 5 = 25 + 5 = 30

OpenStudy (anonymous):

ok...cool thanx